OpenStudy (anonymous):

Solve-6(5-2z)+5z=8z-3(16-3z) Please show all work

OpenStudy (anonymous):

first distribute the -6 into the parentheses on the left side. Next distribute the -3 into the parentheses on the right side.

OpenStudy (anonymous):

Then combine like terms

OpenStudy (anonymous):

Finally isolate and solve for z

jhonyy9 (jhonyy9):

-30 +12z +5z = 8z -48 +9z 18= -17z +17z 18=0 - so i think you have missed anything or you have wrote one sign wrong - check it please !

OpenStudy (anonymous):

It wrote it right

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=6%285-2z%29%2B5z%3D8z-3%2816-3z%29 Think it is not a -6 at the beginning, it's just a dash...